Solve if possible or determine no solution or infinite solutions.
x+3y=-15
2x+5y=-21

Answer by checkley71(8403)

X=-3Y-15 THEN 2(-3Y-15)+5Y=-21 OR -6Y -30+5Y=-21 OR -Y= 9 OR Y=-9
THEN X+3(-9)=-15 OR X-27=-15 OR X=12